OpenStudy (anonymous):

i need help with this problem because i cant look the answer up in the back of the text book

OpenStudy (anonymous):

yeah go ahead dear friend

OpenStudy (anonymous):

for christmas steve bought 5 more cds than dvds as presents for his freinds. he bought 21 in all . how many cds did he buy?

OpenStudy (anonymous):

means number of cd together with dvd is equal to ---------how much

OpenStudy (anonymous):

If you write these as equations, it helps in the analysis. CD=DVD+5 CD+DVD=21 CD=21-DVD 21-DVD=DVD+5 16=2DVD DVD=8 Which means CD=13. We then check if the solution satisfies the original equation. 8+13=21 and there are 5 more CDs than DVDs :)

OpenStudy (anonymous):

So CDs = DVDs + 5 CDs + DVDs = 21 Substitute. DVDS + 5 + DVDs = 21 2DVDs = 16 DVDs = 8 CDs = 13

OpenStudy (anonymous):

onn the paper ir says that step1. define veriable step2. right question step3.solve step4.wright concution sentence

OpenStudy (anonymous):

so u have variable CD & DVD

OpenStudy (anonymous):

and u purchased 5 CD's more than the DVD so No. of CD = No. of DVD + 5

OpenStudy (anonymous):

Now total no is equal to 21

OpenStudy (anonymous):

so on my paper i should have x=x+5

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now what is x = it is number of dvds

OpenStudy (anonymous):

Not x=x+5, x=y+5 is more appropriate.

OpenStudy (anonymous):

yes dear helper cd is more No. of CD = No. of DVD + 5

OpenStudy (anonymous):

ok that was a good tip polrek

OpenStudy (anonymous):

now i get it

OpenStudy (anonymous):

we can keep No. of CD = x +5, No. of DVD x

OpenStudy (anonymous):

x = x - 5 -x -x -------- -5 whould be the answer

OpenStudy (anonymous):

no no we have to add number of cd with no of dvd

OpenStudy (anonymous):

i.e., x + x + 5 = ?

OpenStudy (anonymous):

7x

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

no 5 can't be added into x because it dint have x

OpenStudy (anonymous):

2x+5

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

which is equal to?

OpenStudy (anonymous):

what do you mean to equal

OpenStudy (anonymous):

means total i.e., no. of cd + no. of dvd = ?

OpenStudy (anonymous):

read question

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

5 divided by 2

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

no dear 2x+5 = 21

OpenStudy (anonymous):

now solve it

OpenStudy (anonymous):

i forgot the 21 i didnt see that sorry

OpenStudy (anonymous):

answer is 16

OpenStudy (anonymous):

now tell me x is equal to ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

no look at the equation carefully

OpenStudy (anonymous):

2x = 21 - 5 => 2x = 16 => x = 16 /2 => x = ??

OpenStudy (anonymous):

answer is 8

OpenStudy (anonymous):

gr8 u did it now u got to know how to reach this point

OpenStudy (anonymous):

but x is number of dvd so number of cd will be equal to ??

OpenStudy (anonymous):

i feel like an idiot because it was so obvios but i got it right finally thank you very mutch

OpenStudy (anonymous):

13 cds

OpenStudy (anonymous):

no of cd = no of dvd+ 5 no of cd = x+ 5

OpenStudy (anonymous):

i want to give u another medal but i cant not possible here

OpenStudy (anonymous):

if u having any other question then share........

OpenStudy (anonymous):

i wish i could give you a real well the one with the the rectangle i still ned help with

OpenStudy (anonymous):

yeah go ahead by the way whts the time there

OpenStudy (anonymous):

i have 408=134+1x

OpenStudy (anonymous):

5:50 im in virginia

OpenStudy (anonymous):

ok that one perimeter = 2 * (L+W) => 408 = 2*(L+W)

OpenStudy (anonymous):

i had seen a movie that was from virginia...... A Walk To Remember...................ultimate one

OpenStudy (anonymous):

408 = 134 + X -134 -134 ------------- 274=1x --- --- 1 1 x=274

OpenStudy (anonymous):

solve the equation...b4 i complete my cigarrette give me answer 408 = 2*(L+W) L is given in the question

OpenStudy (anonymous):

408 = 2 x (L+W)

OpenStudy (anonymous):

L

OpenStudy (anonymous):

=134

OpenStudy (anonymous):

134X274=36,716cm sqaurd

OpenStudy (anonymous):

yes L = 134 L+W = 408/2 L+W = ??

OpenStudy (anonymous):

134+274=??? what does the /2 mean

OpenStudy (anonymous):

does that mean sqaurd

OpenStudy (anonymous):

means 408 divided by 2 =

OpenStudy (anonymous):

204

OpenStudy (anonymous):

then 204 = 134 + w

OpenStudy (anonymous):

oh 204-134=70 70+134=204

OpenStudy (anonymous):

gr8 so width is equal to ??

OpenStudy (anonymous):

274

OpenStudy (anonymous):

no

OpenStudy (anonymous):

204

OpenStudy (anonymous):

no 70 is the width

OpenStudy (anonymous):

yes u r right. 70 cm is the width hey can i ask u sumthing if u dont mind

OpenStudy (anonymous):

so 134 x 70

OpenStudy (anonymous):

you can ask

OpenStudy (anonymous):

no it will give u the area of the rectangle........ whts ur good name and how old u r?

OpenStudy (anonymous):

joey pedro 14 9th grade

OpenStudy (anonymous):

nice name

OpenStudy (anonymous):

yeah you whant the area of the rectangle the question says that you to find the area

OpenStudy (anonymous):

yes then area is equal to 134*70 = 9380 cm^2

OpenStudy (anonymous):

yes finally.

OpenStudy (anonymous):

next question

OpenStudy (anonymous):

are you a teacher ar a student? you should be a teacher

OpenStudy (anonymous):

i have too go dinner time

OpenStudy (anonymous):

yeah u got it ryt.....but i dont teach school kids i teach old post graduate fellows